caira12:

When solving -1/8 (x + 35) = −7, what is the correct sequence of operations? (5 points) Select one: a. Multiply each side by -1/8 , add 35 to each side b. Multiply each side by -1/8 , subtract 35 from each side c. Multiply each side by −8, subtract 35 from each side d. Multiply each side by −8, add 35 to each side

caira12:

?

Logic007:

Step 1 :Multiply both sides by -8

Logic007:

\(-8\times-\frac{1}{8}(x+35)=-7\times-\frac{1}{8}\)

Logic007:

\(x+35=\frac{7}{8}\)

Logic007:

Step 2: Subtract both sides by 35 and solve it

Logic007:

\(x+35-35=\frac{7}{8}-35\)

caira12:

so the answer is option d

Logic007:

Hint : look at my steps

caira12:

so its not option d

Logic007:

Can u tell me why the answer is D?

caira12:

ohhh its option C not d

Logic007:

Yep,u got it right

caira12:

because in step one you multiplied both sides by -8

Logic007:

yep ;) then,step 2,u subtract both sides by 35

caira12:

thank you sm

Logic007:

np ^